OpenStudy (anonymous):

Using complete sentences, explain how to find the equation of the line, in standard form and slope-intercept form, passing through (-1, -7) and (1, -1). Also compare the benefits of writing an equation in standard form to the benefits of writing an equation in slope-intercept form.

OpenStudy (gorv):

given the points......(x1,y1) and (x2,y2) slope = (y2-y1)/(x2-x1)=(-1-(-7))/(1-(-1))=(-1+7)/(1+1)=6/2=3

OpenStudy (gorv):

standard form of line passing through a point & with slope m will be y-y1=m*(x-x1) or y-y2=m*(x-x2)

OpenStudy (gorv):

y-(-1)=3*(x-1) y+1=3x-3 y=3x-3+1 y=3x-2 where y intercept is -2